OpenStudy (anonymous):

Find the equation of tangent to the curve y = ln x at the point where x = a. Hence, find the equation of tangent to the curve which passes through the origin. Given that line y = mx intersects the curve the curve y = ln x at two distinct pts, write down an inequality for m.

OpenStudy (amistre64):

what do we get for the derivative?

OpenStudy (amistre64):

the derivative defines the slope at any given value of x; therfore the tangent line at any given points is defined as: y = f'(a) (x-a) + f(a)

OpenStudy (amistre64):

then your next task seems to define the range of value for m such that lnx = mx has only 2 solutions

OpenStudy (anonymous):

Yes and then???

OpenStudy (amistre64):

and then you come up with the results ...

OpenStudy (anonymous):

I still dont get the first part.

OpenStudy (amistre64):

then ask questions about it; what dont you understand?

OpenStudy (anonymous):

I differentiate ln x i get 1/x

OpenStudy (amistre64):

good, and 1/x at any given x=a is just 1/a

OpenStudy (anonymous):

Correct

OpenStudy (amistre64):

y = 1/a (x-a) + ln(a) is the slope of the tangent line

OpenStudy (amistre64):

lol, equation of the tangent line

OpenStudy (anonymous):

*x-a + ln a * i see

OpenStudy (amistre64):

when does this equation have an x and y intercept at the origin?

OpenStudy (anonymous):

Lol i only know that method

OpenStudy (amistre64):

in other words, for what value of a is (0,0) a good solution to it?

OpenStudy (amistre64):

y = 1/a (x-a) + ln(a) 0 = 1/a (0-a) + ln(a) 0 = -1 + ln(a) 1 = ln(a) , when a=e

OpenStudy (amistre64):